Air is a mixture of gases with the following mass percentage composition: 75.52% N2, 23.15% O2, 1.28% Ar, and 0.046% CO2 (a) What are the partial pressures of N2, O2, Ar, and CO2, when the total pressure is 1.100 atm? (b) Calculate the molar Gibbs energy of mixing of air at 25 °C assuming ideal-gas behavior. (c) Determine the molar enthalpy of mixing and the molar entropy of mixing for air at 1.100 atm and 25 °C.5. (3 pts) Air is a mixture of gases with the following mass percentage composition: 75.52% N2, 23.15% O2, 1.28% Ar, and 0.04
Answer #1

by 75.52 Ang = mole fraction of N2 = 75.52f 23:15 + 1-28+0.046 = 0·7552 NO2 = mole fraction of o2 = 23.15 76-52723.151.11% +